OpenStudy (anonymous):

(y-1)/(y+1)= c((x-1)/(x+1)) how would you solve for y here?

OpenStudy (anonymous):

(y - 1)(x + 1) = c(x - 1)(y + 1) yx + y - x - 1 = cyx + cx - cy - c y(x + 1) - x - 1 - cyx + cy = cx - c y(x + 1 + c - cx) = x(c + 1) - c + 1 y(x(1 - c) + 1 + c) = x(c + 1) - c + 1 y = [x(c+1) - c + 1]/[x(1-c) + 1 + c]

OpenStudy (anonymous):

y cannot be -1, nor can x. the simplest case is when c = 1, in which case y = x

OpenStudy (anonymous):

thanks, brother